Hi , i'm trying to integrate 1/(2+6x+9x^2). I know we're supposed to complete the square as part of our steps and I got 1/(3x+3)^2-7. I'm not sure if i'm correct with what I did and i'm also stuck from there. Your help would be appreciated.
Answer #1

2 +63 +94? (32)2+2.38.1+1 +1+1 Gu+1)2+1 I= S H an 2 +67+974 da s (3u+1)*+1 = 34+1 = tano zdn= See odo sec od tanot 3 Sec²o do